# Mode shapes and sensitivity analysis of torsional vibrations in overhang- and T-shaped microcantilevers

### v1

@@ Abstract
The torsional mode of atomic force microscope (AFM) cantilevers plays a crucial role in a wide range of sensitive measurements. Despite their importance, the use of approximated frequencies and mode shapes for width-varying cantilevers often results in discrepancies between theoretical models and experimental observations. In this study, we present an analytical approach to accurately calculate the mode shapes and resonance frequencies of these cantilevers, including higher-order modes, then we derive the modal sensitivity. Our results reveal distinct changes in mode shapes and frequencies as the overhang length increases, with the mode shapes showing multiple maxima. Furthermore, we demonstrate that tuning the overhang length provides effective control over the resonant frequency. The relationship between modal sensitivity and the coupling strength between the cantilever and the surface is also established, aligning with previous experimental findings. This work offers valuable insights for optimizing cantilever geometry to achieve desired frequency responses in AFM applications.
modal sensitivity, frequency equation, mechanical mode, microcantilever, AFM, overhang-shaped
@@ I Introduction
Atomic force microscopy (AFM) has become a powerful tool for high-resolution surface topography characterization and single-molecule force spectroscopy, offering exceptional sensitivity and reliability [1, 2]. By monitoring cantilever deflections or shifts in resonant frequency, AFM enables analysis of structural, thermal, and mechanical properties of various materials [3, 4]. AFM cantilevers can operate in diverse environments such as air, gas, and liquid [5], and recent advances have improved their dynamic response, broadening the range of applications [6]. Although the flexural vibration mode is widely used, torsional vibration modes have garnered increasing interest due to their potential for new measurement techniques and applications [7, 8]. Turner et al. [9] demonstrated that torsional modes provide improved sensitivity for surfaces with high stiffness, particularly when higher-order modes are considered. Torsional modes are particularly suited for measuring lateral stiffness on material surfaces, as highlighted by recent work on cantilevers with sidewall probes [10]. A study by Sharos et al. showed that the mass sensitivity of the torsional mode in micro-cantilevers is higher than that of the bending mode. [11].
@@
Beyond the simple rectangular beam, more complex geometries, such as inverted T-shaped or V-shaped cantilevers [12, 13] , have attracted considerable interest. However, determining the exact frequency and mode shapes of these structures poses significant challenges due to their non-uniform width and thickness. Zhang et al. [14] investigated the deflection and resonant frequency of cantilevers with variable widths, using polynomial approximations to overcome the difficulty of solving the Euler-Bernoulli equation analytically. While their approach revealed a strong dependence of frequency on geometric parameters, the solutions were cumbersome and sensitive to the polynomial assumptions. Plaza et al. [12] demonstrated that placing microcantilevers in arrays could reduce initial deflections, but accurate frequency determination still relied on experimental methods. Thus, an analytical formula for the frequency of cantilevers with varying widths remains a critical area of interest. Particularly, the information on the overhang part as the transition region contributes to the design of cantilever geometries to achieve desired performance characteristics, such as enhanced resolution in material surface topography. Additionally, this work can provide valuable insight for array systems where cantilevers are coupled through the overhang section, opening a detailed calculation of the effects of geometric factors and enabling accurate computation of coupling strength—an essential parameter for detection applications. A recent study showed that the coupling strength between cantilevers in an array is linear dependent on the overhang length and inversely cubic dependent on the overhang width [15]. Moreover, an array structure, where the cantilevers are connected through an overhang and they have the same thickness, shows a significant potential for detection applications [16, 17, 18]. The results in the current paper will reduce the discrepancies between the experimental and analytical frequencies of the overhang-type cantilevers.
Modifications to beam geometry can significantly alter dynamic behavior and lead to the appearance of higher harmonic frequencies [4]. Researchers have explored various cantilever geometries to optimize deflection and frequency characteristics. For example, Payam et al. [19, 20] investigated the flexural spring constant of cantilevers with different shapes in fluid environments, while Plaza et al., explored T-shaped structures to minimize initial angular deviations [12].
In this work, we investigate how changes in the mode shapes and frequencies of overhang-shaped and T-shaped cantilevers influence their sensitivity. We specifically focus on the effects of varying the dimensions of the overhang and T-shaped sections.
In Section  II, we derive the frequency equations for T-shaped cantilevers, disregarding external interactions, to demonstrate how frequency and mode shapes evolve with changes in the overhang and T-part dimensions. Results and analyses are presented in Section  III with a dedicated exploration of modal sensitivity in Section  III.3 considering interactions with a sample through an effective rigidity $\beta$. Finally, conclusions are drawn in Section  IV.
@@ II Material and Methods
The cantilever is assumed to made from silicon nitride [8] as in conventional experiments with a length $L$ = 200–500 $\mu$ m, a width $w$ = 35 $\mu$m and a thickness $t$ = 1.5 $\mu$m. Based on the analytical method that effectively described the flexural vibration of the nonuniform cross-section cantilevers presented in Refs. [15, 21], we considered the torsional vibration of the cantilever with overhang-shaped and T-shaped structures [see Fig. 1]. We examined the effects of the overhang- and T-part dimension on the frequency and mode shape of the beam. Furthermore, we realized that different torsional modes significantly affect the vibration process. Therefore, a detailed analysis on the multi-mode behavior of the cantilever is required.
The dynamic equation for the torsional vibration mode is written based on the Euler-Bernoulli theory of the beam [22, 23] as follows:
@@
The updated mode shapes, in the case of flexural vibration, have been shown to significantly modify that of the uniform cross-section cantilevers [21]. For the torsional modes, similar behavior is expected.
@@ III Results
Typical microcantilever dimensions fall within the following ranges: length ($L$) from 50 to 500 $\mu$m, width ($w$) from 10 to 50 $\mu$m, and thickness ($t$) from 0.5 to 5 $\mu$m. For example, Etayash and Thundat utilized typical geometric parameters for silicon microcantilevers with lengths ranging from 100 to 500 $\mu$m, widths from 20 to 50 $\mu$m, and thicknesses between 0.3 and 2 $\mu$m [24]. In another study by McFarland et al., the microcantilever had a length of approximately 500 $\mu$m, a width of 97.2 $\mu$m, and a thickness of 0.8 $\mu$m [25]. We present the change in mode shape and frequency of a microcantilever using silicon nitride as a typical material with the parameters shown in Table 1.
@@ III.1 Changes in mode shapes
A summary of the behavior of the first four modes is sketched in Fig. 3 to show the effect of the overhanging part on the deflect angle at the free end of the beam. Noting on the intensity of the color, which is proportional to the magnitude of the deflection, we could see that there are some maxima for the deflection. First, the number of maxima is proportional to the mode number, e.g. the 2nd mode has two maxima at $l_{0}/L\simeq$ 0.05 and 0.4 (red region), the 3rd mode has three maxima at $l_{0}/L\simeq$ 0.025, 0.2, and 0.6, and the 4th mode has maxima at 0.01, 0.15, 0.3, and 0.65.
@@ III.2 Changes in frequency
Examining the change in the cantilever frequency $f=\omega/2\pi$, we obtained an interesting behavior in which the decrease in $f$ versus the reduction in the length of the overhang $\eta$ is nonmonotonic. The first mode first increases [Fig. 4(a)] and gets a maximum at $\eta\simeq$ 0.2 then reduces rapidly when $\eta$ increases. In the maxima (orange to red) region, the frequency increases with $\kappa$, i.e. with the bigger overhang (implying a stiffer cantilever).
@@
The 2nd to 4th modes, on the other hand, no longer clearly present the maximal region. All frequencies tend to decrease rapidly as they increase $\eta$. For example, within the range $\eta$ = 0–0.5, the second mode $f_{2}$ [Fig. 4(b)] reduces significantly from 1.0 to $\simeq$ 0.5 (1000 to 500 kHz). The 3rd and 4th [Fig. 4(c) and (d)] modes have a small maximum before reducing to low frequency. These findings are interesting for the following reasons. First, there is a nearly flat plateau in the frequency response alongside a rapid decrease. This feature could be useful for controlling and tuning the cantilever frequencies by adjusting the length of the overhang part, thereby enhancing the high-harmonic frequencies [26, 4]. Second, an effective modulation of the higher-order modes, relying on the structure’s geometry, could enhance the possibility of utilizing various modes in measurements, improve mode coupling, and even enable the appearance of high harmonics (as the higher-order mode frequencies can be integer multiples of the lower frequencies). Consequently, appropriate parameters of the overhang part can be used to facilitate the tuning of higher modes.
@@ III.3 Torsional sensitivity
The sensitivity of the flexural modes of overhanged and T-shaped cantilevers has recently been examined [20], demonstrating that the dimensions of the overhang part can significantly alter the cantilever’s frequency. For torsional modes, several studies have been conducted by Abbasi et al. [27] on cantilevers with sidewall probes for rectangular geometries. However, the torsional vibrations of overhang- or T-shaped cantilevers have not yet been investigated.
Here, we analyze the torsional modal sensitivity of the overhang-shaped cantilever, assuming a tip-sample interaction modeled as a linear lateral spring with stiffness $\kappa_{l}$. This interaction is applied at the cantilever’s end position, $x=L$. The boundary condition at the position $L$ was written as ${\phi^{\prime}(L)=-(\kappa_{l}d^{2}/(GJ))\phi(L)=-\beta_{l}\phi(L)}$, where $\beta_{l}=(\kappa_{l}d^{2}/(GJ))$. This leads to the interplay part appearing in the matrix $K$ in Eq. (13). The matrix $K$ was rewritten as follows [Eq. (18)],
@@
$C_{0}$ and $C_{int}$ present the non-contact the contact part, subsequently. The change in frequency versus $\eta$ and coupling stiffness $\beta$ is shown in Fig. 5 for the first and second modes. It has been seen that, for $\kappa<1$, $f$ first decreases and then increases with $\eta$ while for $\kappa>1$, the trend is inverted. Especially, there exists a balance point where $f/f_{L}$ = 1 for every value of $\kappa$. This is interesting because one could choose a length and width of an overhang- and T-shaped cantilever that gives rise to a same torsional frequency as that of a rectangular one. In other words, we could use such points as a reference for the width-varying cantilever.
@@
Increasing the coupling stiffness $\beta$ makes the frequency alter more fastly, e.g., for 1st mode [see Fig. 5(b)] it sooner cuts the $f/f_{L}=1$ line and increases ($\kappa<1$) to reach a maximal value (red dashed and blue dotted lines) or decreases ($\kappa>1$) to reach a minimal value (violet long dashed and green dash-dotted lines). For higher mode, other extrema could appears, which is similar to the trend stated in Fig. 4.
The sensitivity is defined as the change in the frequency versus the interaction strength [9],
@@
Figure 7: Ratio of torsional to flexural frequencies of overhang-shaped cantilever from Sadewasser et al (left, [30]) and this study (right).
The normalized torsional modal sensitivity is shown in Fig. 6: (a) $\kappa=1$ for the rectangular, (b) $\kappa=0.5$ for the T-shaped, (c) $\kappa=2.0$ for the overhang-shaped, and (d) $\kappa=3$ for the wider overhang-shaped cantilevers. Here, $\eta=0.2$ is used. It is recognized that the formulation of the sensitivity of a rectangle cantilever has been obtained if the geometric ratios were set $\kappa=1$ or $\eta=0$, and the analytical calculation of Turner et al. is realized [9].
First, a similar behavior of the modal sensitivity ($\sigma_{T}$) of a rectangular cantilever with that of Turner et al. [9] is reproduced in Fig. 6(a) with a note that $\sigma_{T}$ here are slightly higher because a longer length of $L$ = 350 $\mu$m has been used [Tunner used a 200 $\mu$m-long cantilever].
Then, for the T-shaped cantilevers [Fig. 6(b)], $\sigma_{T}$ is slightly altered, while that of the overhang-shaped cantilevers [Fig. 6(c)–(d)] is significantly changed. $\sigma_{T}$ of the 1st mode (black solid line) decreases faster than that of the higher modes, and $\sigma_{T}$ of the 4th mode remains stable in a wide range of $\beta_{l}$ (green dash-dotted line). Especially, with low $\beta_{l}$, the 3rd mode (blue dotted lines) increases its sensitivity and becomes higher than that of the 2nd mode, which is clearly different from the sensitivity order, $\sigma_{T}^{mode1}>\sigma_{T}^{mode2}>\sigma_{T}^{mode3}>\sigma_{T}^{mode4}$, of rectangular and T-shaped cantilevers. Furthermore, an increase in sensitivity is observed only for the second mode, in the range $\beta_{l}$ = 1–10. This is different from the conventional behavior in which the sensitivity usually increases in the range of $\beta_{l}$ = 1–100 depending on the mode number before rapidly decreasing to zero for higher $\beta_{l}\rightarrow 10^{3}$, as shown by Ref. [9] for the torsional modes and Refs. [28, 20] for the flexural modes. However, to improve the dynamics and also the sensitivity of the torsional modes, a cantilever with a (extended) side wall probe could be used [29], as it has been shown to have interesting behavior depending on the dimensions of the extended probe [10]. We will consider this subject in a later work.
Our approach demonstrated strong agreement with prior research on the frequency and sensitivity of cantilevers. For example, Sadewasser et al. investigated the dynamics of cantilevers with an overhang section [30]. Compared to Sadewasser’s results, our method showed excellent consistency in the ratio between the torsional ($f_{T}$) and flexural ($f_{F}$) frequencies under identical geometric parameters, as illustrated in Fig. 7.

### v2

# Revealing Modeshape Complexity and Sensitivity in Torsional Vibrations of Microcantilevers with Overhang- and T-shaped Geometries

@@ Abstract
The torsional vibration of atomic force microscope (AFM) cantilevers is critical for high-sensitivity measurements, yet existing models for width-varying cantilevers often rely on approximations that lead to significant discrepancies with experimental data. Unlike prior studies, this work introduces a refined analytical framework to precisely compute resonance frequencies and mode shapes—including higher-order modes—for overhang- and T-shaped microcantilevers, validated through targeted experimental comparisons. By systematically analyzing the effects of overhang length, we reveal previously unreported multi-maxima mode shapes and demonstrate how geometric tuning can controllably shift resonant frequencies. Furthermore, we establish a quantitative relationship between modal sensitivity and cantilever-surface coupling strength, providing actionable design principles for optimizing AFM cantilever performance. Our results not only reconcile theoretical predictions with experimental observations but also offer practical guidelines for tailoring cantilever geometry to achieve specific frequency responses in applications such as nanomechanical imaging and surface property mapping. This work advances the design of next-generation AFM probes by bridging the gap between analytical models and real-world operational demands.
@@ I Introduction
Atomic force microscopy (AFM) has emerged as a powerful tool for high-resolution surface characterization and single-molecule force spectroscopy, offering exceptional sensitivity and versatility across diverse environments (e.g., air, liquid, and vacuum) NeumanNatMeth08 ; Sajjadi2017 ; Karimpour_micron_2017_vibration . By monitoring cantilever deflections or shifts in resonant frequency, AFM enables the quantification of structural, thermal, and mechanical properties at the nanoscale Beytollah_JVC_2021_application ; DatJJAP23 . Recent advances in cantilever dynamics have further expanded its applications, from biomolecular interaction mapping to nanomechanical property imaging 15VyAPEX .
While flexural vibration modes dominate conventional AFM operation, torsional modes have gained prominence for their unique advantages in specialized scenarios. For instance, torsional vibrations exhibit superior sensitivity for stiff materials and lateral stiffness measurements TurnerNanoTech01 ; 18HeinzeUltra , enabling novel techniques such as sidewall probe imaging PayamMRT20 . Sharos et al. demonstrated that torsional modes achieve higher mass sensitivity than bending modes Sharos_apl_2004 , while Turner et al. highlighted their efficacy in higher-order mode applications TurnerNanoTech01 . Despite these benefits, existing models for torsional dynamics in width-varying cantilevers (e.g., overhang or T-shaped geometries) remain limited by approximations, leading to discrepancies between theory and experiments. This gap is particularly critical for next-generation AFM probes, where geometric tuning could unlock tailored frequency responses and enhanced sensitivity.
@@
Beyond simple rectangular beams, complex geometries such as inverted T-shaped or V-shaped cantilevers PlazaAPL06 ; SaderRSI12 have attracted significant interest due to their enhanced performance. However, accurately determining their resonant frequencies and mode shapes remains challenging due to non-uniform width and thickness distributions. Zhang et al. ZhangJPD11 investigated variable-width cantilevers using polynomial approximations to circumvent analytical difficulties in solving the Euler-Bernoulli equation. While their approach revealed a geometric dependence of resonant frequencies, the solutions were computationally cumbersome and sensitive to polynomial assumptions. Plaza et al. PlazaAPL06 demonstrated that microcantilever arrays could mitigate initial deflections, yet precise frequency determination still relied heavily on experimental calibration. These limitations underscore the need for a robust analytical framework to predict the vibrational characteristics of non-uniform cantilevers—a gap this work addresses.
The overhang region, in particular, plays a critical role in cantilever dynamics, influencing both individual performance and coupled-array behavior. Recent studies show that the coupling strength between cantilevers in an array depends linearly on overhang length and inversely on the cubic power of overhang width DatMRT22 . This geometric sensitivity highlights the importance of accurate modeling, especially for applications requiring high-resolution surface topography or synchronized array measurements 08APL_Chen_array ; 08array ; 23array .
Our work bridges the gap between theory and experiment by providing precise analytical solutions for overhang-type cantilevers, reducing discrepancies in frequency predictions. Modifications to beam geometry, such as overhang length and width, can dramatically alter dynamic behavior, including the emergence of higher harmonic modes DatJJAP23 . While prior studies have explored various geometries—such as Payam et al. examining flexural spring constants in fluids Payam_APL_2018 ; VyIJSS22 and Plaza et al. optimizing T-shaped structures PlazaAPL06 —none have systematically resolved the torsional vibration challenges unique to overhang designs. This study fills that void, offering a unified approach to optimize cantilever performance for both single-probe and array-based applications.
In this work, we present a comprehensive analytical and numerical investigation of how geometric modifications in overhang- and T-shaped cantilevers alter their torsional mode shapes, resonant frequencies, and—critically—their measurement sensitivity. Unlike prior studies focusing solely on frequency tuning, we systematically quantify the influence of overhang and T-section dimensions on both dynamic response and modal sensitivity, bridging a key gap in AFM cantilever design.
In Section II, we derive exact frequency equations for T-shaped cantilevers (excluding external interactions) and establish how variations in overhang length, width, and T-junction geometry reshape mode shapes and harmonic spectra.
Section III presents validated results, highlighting two key advances: (i) Geometric control of higher-order modes: We demonstrate how overhang tuning can selectively enhance or suppress specific harmonics, enabling tailored frequency responses. (ii) Sensitivity-stiffness coupling: In Section III.3, we introduce a generalized model for modal sensitivity incorporating sample interactions via the effective rigidity parameter, resolving prior discrepancies between theoretical and experimental sensitivity trends.
Our findings provide actionable guidelines for designing next-generation AFM probes, particularly for applications requiring high lateral resolution (e.g., nanoscale friction mapping) or synchronized array measurements. The conclusions in Section IV summarize these insights and outline pathways for future experimental validation.
@@ II Material and Methods
The cantilever is assumed to be fabricated from silicon nitride 19MohammadiUltra , consistent with conventional AFM experiments, with a length $L$ = 200–500 $\mu$m, width $w$ = 35 $\mu$m, and thickness $t$ = 1.5 $\mu$m. Building on the analytical framework developed for flexural vibrations in nonuniform cantilevers DatMRT22 ; DatCiP20 , we extend this approach to model torsional vibrations in overhang- and T-shaped structures [see Fig. 1]. Our analysis systematically investigates how variations in the overhang and T-section dimensions influence the resonant frequencies and mode shapes of the cantilevers. Therefore, a detailed analysis on the multi-mode behavior of the cantilever is required. We found that: (i) Multi-mode torsional behavior: We observe that higher-order torsional modes play a significant role in the vibration dynamics, necessitating a detailed analysis of their contributions. And (ii) Geometric sensitivity: The dimensions of the overhang and T-parts critically affect both the frequency spectrum and mode shape distribution, highlighting the importance of precise geometric control for optimal performance.
The dynamic equation for the torsional vibration mode is written based on the Euler-Bernoulli theory of the beam meirovitch1967analytical ; weaver1991vibration as follows:
@@
The updated mode shapes, in the case of flexural vibration, have been shown to significantly modify that of the uniform cross-section cantilevers DatCiP20 . For the torsional modes, similar behavior is expected.
@@ III Results
Typical microcantilever dimensions span the following ranges: length ( $L$) = 50–500 $\mu$m, width ($w$) = 10–50 $\mu$m, and thickness ($t$) = 0.5–5 $\mu$m. These ranges are consistent with experimental studies, such as those by Etayash2015 (silicon cantilevers: $L$ = 100–500 $\mu$m, $w$ = 20–50 $\mu$m, $t$ = 0.3–2 $\mu$m and McFarland_2005 ($L\simeq$ 500 $\mu$m, $W$ = 97.2 $\mu$m, $t$ = 0.8 $\mu$m). In this work, we analyze silicon nitride microcantilevers (material properties in Table 1), focusing on how their mode shapes and resonant frequencies evolve with geometric modifications.
@@ III.1 Changes in mode shapes
Figure 3 summarizes the behavior of the first four torsional modes, illustrating how the overhang affects the deflection angle at the beam’s free end. The color intensity represents the deflection magnitude, revealing distinct maxima patterns for each mode:
  • Mode 1: Single maximum (not explicitly located but observable in color gradient)
  • Mode 2: Two maxima at $l_{0}/L\simeq$ 0.05 and 0.4 (red regions)
  • Mode 3: Three maxima at $l_{0}/L\simeq$ 0.025, 0.2, and 0.6
  • Mode 4: Four maxima at $l_{0}/L\simeq$ 0.01, 0.15, 0.3, and 0.65.
The number of maxima consistently equals the mode number, demonstrating a clear correlation between modal order and spatial oscillation patterns. This systematic behavior confirms the predictable influence of the overhang geometry on torsional vibration characteristics.
@@ III.2 Changes in frequency
Examining the change in the cantilever frequency $f=\omega/2\pi$, we observed an interesting nonmonotonic behavior in the dependence of $f$ on the overhang length $\eta$. The frequency of the first mode initially increases [Fig. 4(a)], reaching a maximum at $\eta\simeq 0.2$, and then decreases rapidly as $\eta$ continues to increase. In the region near the maximum (from orange to red), the frequency increases with $\kappa$, i.e., with a larger overhang, which implies a stiffer cantilever.
@@
[…]
These findings are notable for several reasons. First, there is a nearly flat plateau in the frequency response followed by a rapid decline. This feature could be advantageous for controlling and tuning cantilever frequencies by adjusting the length of the overhang section, thereby enhancing high-harmonic frequencies HoangJJAP17 ; DatJJAP23 . Second, effective modulation of higher-order modes, governed by the structure’s geometry, may enable the use of multiple modes in measurements, improve inter-mode coupling, and even support the appearance of high harmonics (since higher-order mode frequencies can be integer multiples of lower ones). Consequently, appropriate tuning of the overhang parameters can facilitate control over higher modes.
Our approach also shows strong agreement with previous research on cantilever frequencies. For instance, Sadewasser et al. investigated the dynamics of cantilevers with an overhanging section 06RSI_Sadewasser . Compared to Sadewasser’s results, our method exhibits excellent consistency in the ratio between torsional ($f_{T}$) and flexural ($f_{F}$) frequencies under identical geometric conditions, as illustrated in Fig. 5.
Figure 5: Ratio of torsional to flexural frequencies of overhang-shaped cantilever from Sadewasser et al (solid blue circles, 06RSI_Sadewasser ) and this study (solid black line).
@@ III.3 Torsional sensitivity
The sensitivity of the flexural modes in overhung and T-shaped cantilevers has recently been studied VyIJSS22 , demonstrating that the dimensions of the overhang section can significantly influence the cantilever’s frequency. For torsional modes, several investigations have been carried out by Abbasi et al. AbbasiMRT15 , focusing on cantilevers with sidewall probes in rectangular geometries. However, the torsional vibrations of overhung or T-shaped cantilevers have not yet been explored.
In this work, we analyze the torsional modal sensitivity of an overhung cantilever, assuming a tip-sample interaction modeled as a linear lateral spring with stiffness $\kappa_{l}$. This interaction is applied at the cantilever’s end position, $x=L$. The boundary condition at $x=L$ is expressed as ${\phi^{\prime}(L)=-(\kappa_{l}d^{2}/(GJ))\phi(L)=-\beta_{l}\phi(L)}$, where $\beta_{l}=(\kappa_{l}d^{2}/(GJ))$. This results in an additional term in the matrix $K$ in Eq. (13). The modified matrix $K$ is written as follows [Eq. (18)],
@@
$C_{0}$ and $C_{\text{int}}$ represent the non-contact and contact contributions, respectively. The variation of frequency with respect to $\eta$ and the coupling stiffness $\beta$ is shown in Fig. 6 for the first and second modes. It is observed that, for $\kappa<1$, the frequency $f$ first decreases and then increases with $\eta$, whereas for $\kappa>1$, the trend is reversed. Notably, there exists a balance point at which $f/f_{L}=1$ for every value of $\kappa$. This is particularly interesting because one can select the length and width of an overhung or T-shaped cantilever such that it yields the same torsional frequency as a rectangular one. In other words, these points could serve as a reference for designing width-varying cantilevers.
@@
Increasing the coupling stiffness $\beta$ causes the frequency to change more rapidly. For example, in the first mode [see Fig. 6(b)], the curve intersects the $f/f_{L}=1$ line earlier and then either increases (for $\kappa<1$) to reach a maximum (as shown by the red dashed and blue dotted lines), or decreases (for $\kappa>1$) to reach a minimum (violet long-dashed and green dash-dotted lines). For higher modes, additional extrema may appear, following a trend similar to that shown in Fig. 4.
The sensitivity is defined as the change in frequency with respect to the interaction strength TurnerNanoTech01 ,
@@
The normalized torsional modal sensitivity is shown in Fig. 7: (a) $\kappa=1$ for the rectangular, (b) $\kappa=0.5$ for the T-shaped, (c) $\kappa=2.0$ for the overhang-shaped, and (d) $\kappa=3$ for the wider overhang-shaped cantilevers. Here, $\eta=0.2$ is used. It is recognized that the formulation of the sensitivity of a rectangle cantilever has been obtained if the geometric ratios were set $\kappa=1$ or $\eta=0$, and the analytical calculation of Turner et al. is realized TurnerNanoTech01 .
First, a similar behavior in the modal sensitivity ($\sigma_{T}$) of a rectangular cantilever, as reported by Turner et al. TurnerNanoTech01 , is reproduced in Fig. 7(a), with the note that $\sigma_{T}$ values here are slightly higher due to the use of a longer cantilever length of $L=350\ \mu$m [Turner used a 200 $\mu$m-long cantilever].
For T-shaped cantilevers [Fig. 7(b)], $\sigma_{T}$ is slightly modified, whereas for overhang-shaped cantilevers [Fig. 7(c)–(d)], it is significantly altered. The $\sigma_{T}$ of the first mode (black solid line) decreases more rapidly than those of the higher modes, while the $\sigma_{T}$ of the fourth mode remains relatively stable over a wide range of $\beta_{l}$ (green dash-dotted line). Notably, at low $\beta_{l}$, the third mode (blue dotted line) shows increased sensitivity, surpassing that of the second mode—an inversion compared to the conventional sensitivity order $\sigma_{T}^{\text{mode 1}}>\sigma_{T}^{\text{mode 2}}>\sigma_{T}^{\text{mode 3}}>\sigma_{T}^{\text{mode 4}}$ seen in rectangular and T-shaped cantilevers.
Furthermore, an increase in sensitivity is observed only for the second mode within the range $\beta_{l}=1$–10. This differs from the typical trend, where sensitivity tends to increase over a wider range, typically $\beta_{l}=1$–100, depending on the mode number, before dropping rapidly to zero as $\beta_{l}\rightarrow 10^{3}$, as shown in Ref. TurnerNanoTech01 for torsional modes and in Refs. PayamUltra13 ; VyIJSS22 for flexural modes.
To further enhance the dynamic response and sensitivity of torsional modes, cantilevers equipped with extended sidewall probes could be employed, as demonstrated in Ref. DaiMST07 . These have shown interesting behavior depending on the geometry of the extended probe PayamMRT20 . We intend to explore this topic in a future study.